FEL Upgrade Project Weekly Brief

July 17-21, 2000


The IR Demo was operated in a third harmonic mode generating 100 watt cw at 1 micron this past week. This makes the IR Demo the highest power sub- picosecond laser operating at this wavelength which is very useful for laser material processing and scientific research.


The cost performance report and technical summary for the first month of the project (June 2000) was completed and sent to the ONR and DOE contract monitors.

We received a number of responsive bids for an RFP to provide engineering services for the design and construction of FEL upgrade hardware. The bids are under evaluation according to advertised selection criteria.

Design/Construction Activities

WBS 2 (Facilities)

We made a revised estimate of the total amount of low conductivity cooling water (LCW) that we would need in the FEL facility to accommodate the upgrade. The new estimate (610 gpm for the upgrade and the possible addition of the IBM synchrotron) is comfortably within the presently installed capacity (680 gpm) in the building.

WBS 3 (Beam Physics)

This week we continued work for documentation for the upcoming FEL and Linac conferences. We have completed a preliminary CSR study for the upgrade based on 1-D longitudinal wake field:

- performance at IR Upgrade parameters (160 MeV, 20 mm-mrad, 135 pC) adequate: slight emittance degradation with a small phase space distortion

- performance at "worst case" parameters (100 MeV, 5 mm-mrad, 540 pC) severely degraded, doubled emittance, significant phase space distortion

- bottom line:IR Upgrade numbers seem okay; UV numbers may prove "amusing"

WBS 4 (Injector)

All three spare ceramics for the new high voltage assembly have been ion implanted at LBL. The achieved resistances are in the desired range. A slow increase in the resistance value is noted in measurements taken immediately after implantation and after a few days of air exposure. While the ceramics are at LBL this effect will quantified for the record.

We are continuing our plan for determining the cause for the anomalous high pressure during bake out of the apertured cesiator. No solution thus far. We are suspecting an air leak that opens up in the assembly only during bakeout.



We completed the fabrication of two 7 cell cavities that will be shipped to KEK (Japan) for test of KEK's electropolishing procedure. One cavity has completed chemistry and has achieved 14 MV/m during cold tests. The second cavity has completed chemistry and is installed in the VTA dewar for testing next week.

WBS 9 (Transport)

We are continuing our plan for determining the cause for the anomalous high pressure during bake out. No solution thus far. Beam Transport System (WBS 9) Dipoles* We continued layout of the GX Prototype, verifying with a vendor that our proposal for splitting the magnet at a pole face is the best method of obtaining the tight gap tolerances. Other conceptual effort centered on coil support and connections

* We are still working on characterizing the prototype steel for an order.

* We continue constructing the 3D model of the GX dipole. Quadrupoles

* Pre finish machining of the prototype 3-inch quadrupole yoke parts is due to be completed at the end of next week.

* The winding fixture for the coils is complete. General:

* We are attempting to keep a small part of the time of the two engineers diverted to other duties. We want them available for FEL ongoing projects so that there are minimal gaps in development.

Wiggler (WBS 10.)

* We released a drawing for the core modifications for the optical klystron for the shop. Design of the diagnostics, vacuum pipe, and water manifolds continues. We are nearing completion of the dispersion section details with concentration on instrumentation, wiring connections, covers, and support frame.

Optics (WBS 11)

We started preliminary design this week of a test stand that will be used to test the mirror mounting, figure control and diagnostic configuration.



IR Demo operations went relatively smoothly for this entire week of user operations. During the first part of the week light was delivered to the UVA chemical dynamics collaboration (M. Dewitt and B. Pate) for studies of vibrational energy transfer. During the second half of the week the UVA/ODU atomic physics collaboration (R. Jones, C. Sukenik, et al.) was here to continue their work on harmonic generation in gases and a first attempt at looking the spectroscopy of laser trapped barium atoms. We were able to supply both a higher brightness and better stabilized beam for these experiments compared to their first run back in the March user run. Unfortunately, the atomic group had problems with their atomic sources which terminated the run late Thursday evening. We have used Friday for preparing for next week's operations and to start the vacuum installation for next week's new configuration for detecting Compton x-rays.

All information for prospective users of the FEL is now posted on the JLab web site at: http://www.jlab.org/exp_prog/experiments/index.html